# High Resolution UDF Meshing via Iterative Networks

### v1

@@ A.3 Training convergence
In the method section of the main paper we state that our pipeline applies a sigmoid function to the network outputs before using them as input for the next iteration. We have experimentally found that, when using an identity activation (i.e. direct input) and training with only one iteration, the network goes from an IC of 87.0 on ShapeNet [3] cars at resolution 512 to 88.6, however it starts diverging after that. Using a softmax activation, which helps normalizing the otherwise unbound network outputs, the training converges, but the network does not produce significant improvements over iterations, going from an IC of 87.3 to 87.5 after two iterations. Using a sigmoid activation, instead, showed a more steady improvement over iterations, going from 87.6 to 88.5 after one iteration, and 88.9 after an additional one. However, training it with more than one iteration did not show significant improvements. Using a random number of training iterations, instead, helped the network to converge until 5 iterations, achieving similar IC scores but better CD scores ($9.90\times 10^{-5}$ vs $8.84\times 10^{-5}$), as shown in the main paper, signifying an overall similar accuracy but better surface retrieval capabilities at high resolutions.
A.4 Autodecoder training
For our experiments we used the traditional autodecoder architecture proposed in DeepSDF [24] and used also in [31, 17]. The input meshes are rescaled and centered within a $1,1]^{3}$ volume, and during the data preparation phase, training points are sampled. For each mesh, 20k points are uniformly sampled within the volume, while 400k points are sampled near the surface. To obtain the surface points, 200k points are first uniformly distributed on the surface, and then small amounts of Gaussian noise are added. Gaussian noise with a mean of 0 and a standard deviation of $\sqrt{0.005}$ is applied to the first 200k surface points, and noise with a mean of 0 and a standard deviation of $\sqrt{0.0005}$ is added to the remaining 200k points. The autodecoder network consists of 12 layers, each with 1024 hidden nodes and ReLU activations, and latent codes of size 512. It is trained using L1 loss, without regularization or Fourier encoding, for 10k epochs with a batch size of 16. To focus the network’s capacity on the surface, the UDF is clamped to 0.1. The Adam optimizer [14] is used with learning rates of 0.0005 for the model and 0.001 for the latent codes, with learning rate decay applied at epochs 1600 and 3500 by a factor of 0.35.
A.5 Different UDF architecture
Table A.4: Triangulating a Softplus-based Neural UDF. L2 Chamfer Distance $\times 10^{-5}$ with 2M sample points (CD), F1 score (F1) and Image Consistency (IC) are reported at varying grid resolutions on the ShapeNet [3] cars dataset. The best results are in bold. DCUDF failed to mesh some of the shapes, but its numbers are reported nonetheless.
@@ A.5 Different UDF architecture
To test the robustness of our method to different UDF architectures, we trained an autodecoder with a different architecture: using a softplus activation function and Eikonal loss with a weight of 0.1, with the rest as in the section above. We show the results on ShapeNet [3] cars in Tab. A.4, along with the highest-scoring baselines from the main experiment. While all methods achieved slightly better performance compared to the ReLU-based architecture, the same conclusions apply. Our method outperforms the tested baselines, particularly so at high resolutions, while also achieving lower standard deviations across the dataset.
@@ A.6 Additional figures
We show here additional qualitative results of our method compared to the baselines. In Fig.˜A.1 we show an additional example at different resolutions compared to the NSD-UDF [31] baseline: the shapes look similar at low resolutions, but at 256 and 512 the baseline cannot retrieve large portions of the surface.
@@
Figure A.1: Meshing at different resolutions, additional examples. While NSD-UDF [31] retrieves most of the surface well at a low resolution, it struggles at higher ones. In contrast, our method, recovers the surface well at all resolutions. We use Marching Cubes with both methods.
GT CAP-UDF MeshUDF DCUDF-T NSD-UDF Ours UNDC DMUDF-T NSD-UDF Ours
@@
Figure A.2: Qualitative comparison at resolution 256. Surface meshing results of neural UDFs with all methods at resolution of 256 (and 128 for MGN). indicates that the method is combined with DMUDF. The shapes are the same as in the main text.
GT CAP-UDF MeshUDF DCUDF-T NSD-UDF Ours UNDC DMUDF-T NSD-UDF Ours
@@
Figure A.3: Qualitative comparison at resolution 128. Surface meshing results of neural UDFs with all methods at resolution of 128 (and 64 for MGN). indicates that the method is combined with DMUDF. The shapes are the same as in the main text.
GT CAP-UDF MeshUDF DCUDF-T NSD-UDF Ours UNDC DMUDF-T NSD-UDF Ours
@@
Figure A.4: Additional qualitative comparison at resolution 512. Surface meshing results of neural UDFs with all methods at resolution of 512 (and 256 for MGN). UNDC failed at resolution 512 due to high GPU memory requirements. indicates that the method is combined with DMUDF.
GT CAP-UDF MeshUDF DCUDF-T NSD-UDF Ours UNDC DMUDF-T NSD-UDF Ours
@@
Figure A.5: Additional qualitative comparison at resolution 256. Surface meshing results of neural UDFs with all methods at resolution of 256 (and 128 for MGN). indicates that the method is combined with DMUDF.
GT CAP-UDF MeshUDF DCUDF-T NSD-UDF Ours UNDC DMUDF-T NSD-UDF Ours
Figure A.6: Additional qualitative comparison at resolution 128. Surface meshing results of neural UDFs with all methods at resolution of 128 (and 64 for MGN). indicates that the method is combined with DMUDF.

### v2

@@ A.3 Training convergence
In the method section of the main paper we state that our pipeline applies a sigmoid function to the network outputs before using them as input for the next iteration. We have experimentally found that, when using an identity activation (i.e. direct input) and training with only one iteration, the network goes from an IC of 87.0 on ShapeNet [Chang15] cars at resolution 512 to 88.6, however it starts diverging after that. Using a softmax activation, which helps normalizing the otherwise unbound network outputs, the training converges, but the network does not produce significant improvements over iterations, going from an IC of 87.3 to 87.5 after two iterations. Using a sigmoid activation, instead, showed a more steady improvement over iterations, going from 87.6 to 88.5 after one iteration, and 88.9 after an additional one. However, training it with more than one iteration did not show significant improvements. Using a random number of training iterations, instead, helped the network to converge until 5 iterations, achieving similar IC scores but better CD scores ($9.90\times 10^{-5}$ vs $8.84\times 10^{-5}$), as shown in the main paper, signifying an overall similar accuracy but better surface retrieval capabilities at high resolutions.
A.4 Auto-decoder training
For our auto-decoder experiments we used the traditional auto-decoder architecture proposed in DeepSDF [Park19c]. The input meshes are rescaled and centered within a $1,1]^{3}$ volume, and during the data preparation phase, training points are sampled. For each mesh, 20k points are uniformly sampled within the volume, while 400k points are sampled near the surface. To obtain the surface points, 200k points are first uniformly distributed on the surface, and then small amounts of Gaussian noise are added. Gaussian noise with a mean of 0 and a standard deviation of $\sqrt{0.005}$ is applied to the first 200k surface points, and noise with a mean of 0 and a standard deviation of $\sqrt{0.0005}$ is added to the remaining 200k points. The auto-decoder network consists of 12 layers, each with 1024 hidden nodes and ReLU activations, and latent codes of size 512. It is trained using L1 loss, without regularization or Fourier encoding, for 10k epochs with a batch size of 16. To focus the network’s capacity on the surface, the UDF is clamped to 0.1. The Adam optimizer [Kingma15a] is used with learning rates of 0.0005 for the model and 0.001 for the latent codes, with learning rate decay applied at epochs 1600 and 3500 by a factor of 0.35.
A.5 Different auto-decoder UDF architecture
Table A.7: Triangulating a Softplus-based auto-decoder. L2 Chamfer Distance $\times 10^{-5}$ with 2M sample points (CD), F1 score (F1) and Image Consistency (IC) are reported at varying grid resolutions on the ShapeNet [Chang15] cars dataset. The best results are in bold.
@@ A.5 Different auto-decoder UDF architecture
To test the robustness of our method to different UDF architectures, we trained an auto-decoder with a different architecture: using a softplus activation function and Eikonal loss with a weight of 0.1, with the rest as in the section above. We show the results on ShapeNet [Chang15] cars in Tab. A.7, along with the highest-scoring baselines from the main experiment. While all methods achieved slightly better performance compared to the ReLU-based architecture, the same conclusions apply. Our method outperforms the tested baselines, particularly so at high resolutions, while also achieving lower standard deviations across the dataset.
@@ A.6 Additional figures
We show here additional qualitative results of our method compared to the baselines. In Fig.˜A.1 we show an additional example at different resolutions compared to the NSD-UDF [Stella24] baseline: the shapes look similar at low resolutions, but at 256 and 512 the baseline cannot retrieve large portions of the surface.
@@
In Fig. A.7, we show the meshing on the CAP-L 3D scene [Zhou24] used in Fig. 4 of the main text in greater size to better appreciate the details, e.g., the base of the statues.
@@
Figure A.1: Meshing at different resolutions, additional examples. While NSD-UDF [Stella24] retrieves most of the surface well at a low resolution, it struggles at higher ones. In contrast, our method, recovers the surface well at all resolutions. We use Marching Cubes with both methods.
GT CAP-UDF MeshUDF DCUDF-T DCUDF-nocut NSD-UDF Ours UNDC DMUDF-T NSD-UDF Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure A.2: Qualitative comparison at resolution 256. Surface meshing results of auto-decoder-based neural UDFs with all methods at resolution of 256 (and 128 for MGN). indicates that the method is combined with DMUDF. The shapes are the same as in the main text.
GT CAP-UDF MeshUDF DCUDF-T DCUDF-nocut NSD-UDF Ours UNDC DMUDF-T NSD-UDF Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure A.3: Qualitative comparison at resolution 128. Surface meshing results of auto-decoder-based neural UDFs with all methods at resolution of 128 (and 64 for MGN). indicates that the method is combined with DMUDF. The shapes are the same as in the main text.
GT CAP-UDF MeshUDF DCUDF-T DCUDF-nocut NSD-UDF Ours UNDC DMUDF-T NSD-UDF Ours
@@
Figure A.4: Additional qualitative comparison at resolution 512. Surface meshing results of auto-decoder-based neural UDFs with all methods at resolution of 512 (and 256 for MGN). UNDC failed at resolution 512 due to high GPU memory requirements. indicates that the method is combined with DMUDF.
GT CAP-UDF MeshUDF DCUDF-T DCUDF-nocut NSD-UDF Ours UNDC DMUDF-T NSD-UDF Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure A.5: Additional qualitative comparison at resolution 256. Surface meshing results of auto-decoder-based neural UDFs with all methods at resolution of 256 (and 128 for MGN). indicates that the method is combined with DMUDF.
GT CAP-UDF MeshUDF DCUDF-T DCUDF-nocut NSD-UDF Ours UNDC DMUDF-T NSD-UDF Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure A.6: Additional qualitative comparison at resolution 128. Surface meshing results of auto-decoder-based neural UDFs with all methods at resolution of 128 (and 64 for MGN). indicates that the method is combined with DMUDF.
GT CAP-UDF MeshUDF DCUDF-T NSD-UDF
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours DMUDF NSD-UDF Ours
Refer to caption Refer to caption Refer to caption Refer to caption
Figure A.7: Meshing of a 3D scene from CAP-L [Zhou24]. Surface meshing results of the "Burghers" scene with all methods at resolution of 512. UNDC failed at resolution 512 due to high GPU memory requirements. indicates that the method is Dual Contouring-based, otherwise it is Marching Cubes-based.
A.7 DoubleCoverUDF [Hou23a] tuning and min-cut
To achieve better results with DCUDF [Hou23a] compared to its default parameters, we have run the algorithm 3 times per experiment and per resolution, with different parameters, and we have selected the best run in each scenario. Additionally, in Fig. A.8, we present the meshing results of DCUDF [Hou23a] on UDFs from different models on cars, both with and without the final min-cut step, at resolutions of 128 and 512. Without the min-cut, more surfaces are retained, though they are double-layered. Moreover, the overall meshing quality deteriorates at higher resolutions, following the trend observed for the other baselines.
Auto-decoder cars CAP-L cars [Zhou24] DiffUDF cars [Fainstein24]
GT Refer to caption Refer to caption Refer to caption
128
with min-cut
Refer to caption Refer to caption Refer to caption
without min-cut
Refer to caption Refer to caption Refer to caption
512
with min-cut
Refer to caption Refer to caption Refer to caption
without min-cut
Refer to caption Refer to caption Refer to caption
Figure A.8: DoubleCoverUDF min-cut. Comparing reconstructions of DCUDF [Hou23a] with and without the min-cut step at resolution 128 (middle) and 512 (bottom).
A.8 Meshing ground-truth UDF
Although our method is designed to handle imperfect UDFs, we also evaluate it on true UDFs computed directly from ground-truth meshes to verify that it does not introduce unwanted artifacts. We report results on the ShapeNet [Chang15] cars dataset in Tab. A.8 (top), and visualize several triangulations at a resolution of 512 in Fig. A.9. As observed, our method does not introduce significant artifacts; however, as expected, multiple iterations provide no benefit in this setting.
In the same table, we also compute the number of holes as surface boundaries, similarly to the description provided by DCUDF [Hou23a], at resolution 512. We report the average. We notice that the original meshes have a large number of boundaries because they contain multiple detailed components and inner structures. None of the methods faithfully recovers the correct mesh topology. Methods that rely directly on MC triangulation, such as CAP-UDF [Zhou24], NSD-UDF+MC [Stella24] and Ours+MC, tend to suffer from micro-holes and gaps between some of the faces. MeshUDF uses a heuristic specifically designed to reduce this behavior and connect as many portions of the surface as possible, explaining the lower number of holes. DCUDF-T [Hou23a] starts from an inflated mesh, so it generally tends to have fewer holes. Simple postprocessing steps can be applied to all methods to improve the final mesh quality. We take as an example the first of the ShapeNet Cars (object 100715345ee54d7ae38b52b4ee9d36a3), and we apply Trimesh-based postprocessing (fill small holes, merge close vertices, remove spurious faces), showing that all methods improve.
GT
Refer to caption Refer to caption Refer to caption
NSD-UDF [Stella24]
Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption
Figure A.9: Triangulating ground-truth UDF. Surface meshing results on the ground-truth UDF in comparison to NSD-UDF [Stella24]. Our method does not introduce significant unwanted artifacts.
Table A.8: Triangulating ground-truth UDFs. Results are reported on the ShapeNet [Chang15] cars dataset, using the true UDF computed from the mesh. We compare against NSD-UDF + MC [Stella24] at multiple resolution with the Median L2 Chamfer Distance $\times 10^{-5}$ with 2M sample points (Chamfer-Distance), and against all MC-based methods with the average number of holes, as surface boundaries at resolution 512(Number of holes). For the latter, we also report results on the first car of the dataset, before and after post-processing the mesh.
(a) Chamfer-Distance
Res.
Method 128 256 512
NSD-UDF + MC [Stella24] 1.34 0.230 0.0300
Ours + MC at iter. 1 1.36 0.244 0.0294
Ours + MC at iter. 6 1.40 0.231 0.0300
(b) Number of holes at resolution 512
ShapeNet cars [Chang15] Car 1 (before post-process.) Car 1 (post-processed)
GT 783 1552 -
CAP-UDF [Zhou24] 10024 13265 2125
MeshUDF [Guillard22b] 428 997 240
DCUDF-T [Hou23a] 7 6 6
DCUDF-T-nocut [Hou23a] 2.65 2 2
NSD-UDF+MC [Stella24] 8280 8885 1387
Ours+MC 9284 9808 1604